OpenStudy (anonymous):

Simple Question. So I have a vector in rectangular form as (-5, -1). I'm converting to polar form -> (r, thetha)

OpenStudy (anonymous):

So, I set this up: \(\ \huge r=\sqrt{x^2 + y^2} = \sqrt{-25^2 + -1^2} \). I get \(\ \huge -\sqrt{26}. \). Is that correct?

OpenStudy (alexwee123):

should be + or - root 26

OpenStudy (anonymous):

How would I convert the rectangular coordinates to polar coordinates?

OpenStudy (dumbcow):

well you're half-way there :) you got r = sqrt(26) now use the fact that tan(theta) = y/x remember you are in 3rd quadrant

OpenStudy (alexwee123):



OpenStudy (anonymous):

@dumbcow Would it be positive sqrt 26 or negative? If positive, how?